Please help me with the following problem
Solve the equation
1/2x2 -4x +8 = 0
this is what i did (1/2x-2)(x-4) = 0
then using the repeat factor of 0, i got x=4

Answer by stanbon(75887)

Solve the equation
(1/2)x^2 -4x +8 = 0
Factor out (1/2)
(1/2)(x^2 - 8x + 16) = 0
Factor again:
(1/2)(x-4)^2 = 0
Solve for "x":
x = 4 with multiplicity two.
===============================
Cheers,
Stan H.
===============================